A 1500 electrical kettle is used to boil 1.0 L of water which was originally at 10 degrees Celsius it takes 5.0 min to boil what is the efficiency of the kettle
1
Expert's answer
2012-02-14T11:14:36-0500
Q=mc(100-t)+mL
Where t=10 – starting temperature
For water from table:
C=4200
L=22.6*105
M=vp=1000kg/m3*0.001m3=1kg
Q=1*4200*90+1*22.6*100,000=2638000dj -needed energy to boil water.
A=1500*5*60=4500000 work what is done.
K=Q/A=59%